OpenStudy (anonymous):

What are the coordinates of the x-intercepts of the graph of y = x2-x-6 & how to find it !

OpenStudy (anonymous):

@dan815 Can you help me with this one please ?

OpenStudy (amistre64):

what is the value of y for any point that is on the x axis?

OpenStudy (anonymous):

Refer to the attached plot.

OpenStudy (shamim):

X intercept means y=0

OpenStudy (shamim):

y=x^2-x-6 y=x^2-3x+2x-6 y=x(x-3)+2(x-3) y=(x+2)(x-3) Right?

OpenStudy (shamim):

If y=0 then 0=(x+2)(x-3) x=?